a-b=25,a+b=81,8ab(a+1/a)^2=?
Answer by ikleyn(53763)

.

The strategy of solution is in two steps.


First step is to solve the system of the two given equations

    a - b = 25

    a + b = 81.


It is VERY EASY: simply add the equations to get

    2a = 25 + 81 = 106,

which implies  a = 106/2 = 53,  and then  b = 81 - a = 81 - 53 = 28.



The second step is to calculate  (8ab)*(a + 1/a)^2.


It is just an elementary job, which I leave to you.